# What sets the metallicity of Ultra-Faint Dwarfs?

### v1

@@ 4.2 The mass loading factor
Redshift evolution of $\eta$ is negligible in the FIRE and FIRE-2 simulations for $40<z<$ 4 Muratov et al. (2015); Pandya et al. (2021), while Mitchell et al. (2020) find a weak redshift dependence at $z<3$, but no evolution at $z>3$. Therefore, $\eta$ is not expected to evolve significantly with redshift, and low redshift measurements can be compared to our results, which most strongly impact UFDs at $z\gtrsim 8$ when they are expected to be star forming.
@@
We find that on average, $\eta\approx\mathrm{const}$ is required to produce flattening in the luminosity–metallicity relation. Rey et al. (2025), for example, find that the inclusion of radiative feedback in addition to SN feedback in their EDGE simulations reduces $\eta$ and leads to higher metallicity in UFDs, making a better fit to luminosity–metallicity observations. In simulations with radiative feedback, the time-averaged $\eta<100$ is relatively constant in stellar mass for their sample of UFDs, with the $68\%$ interval between a few tenths to several hundred over the cosmological history of a UFD. This agrees with $\eta\approx\rm const$ as an explanation for the flattening of the $\mathrm{[Fe/H]-M_{V}}$ relation in the UFD regime, and broadly with our required range of $\eta_{\rm max}$ values.
Rather than regulating star formation with different modes of ISM heating, the outflow wind model in TNG50 determines wind injection velocity by the local dark matter velocity dispersion, and subject to a minimum velocity requirement, therefore the winds launched at the injection scale reach a well defined maximum that leads to flattening in $\eta$ at this scale (Pillepich et al., 2018; Nelson et al., 2019b). However, as Nelson et al. (2019b) show, the emergent outflows at larger radii do not necessarily show the same flattening, subject to the radial velocity and distance used to define galactic outflow.
@@ 5 Summary and Conclusions
    We examine the metallicity distribution of the IGM gas that would be accretable by UFD progenitor galaxies at $105<z<$ 10 in several state-of-the-art simulations of galaxy formation with different treatments of the gas thermodynamics and radiative transfer, star formation, Pop III modelling, and stellar feedback. We find that although the distributions are fairly broad, most of the available IGM gas has low metallicity with only a few per cent of the IGM available to UFDs enriched to $\mathrm{[Fe/H]}\geq-4$ (see Figure 2).

### v2

@@ 4.2 The mass loading factor
Redshift evolution of $\eta$ is negligible in the FIRE-1 and FIRE-2 simulations for $0<z<4$ Muratov et al. (2015); Pandya et al. (2021), while Mitchell et al. (2020) find a weak redshift dependence at $z<3$, but no evolution at $z>3$. Therefore, $\eta$ is not expected to evolve significantly with redshift, and low redshift measurements can be compared to our results, which most strongly impact UFDs at $z\gtrsim 8$ when they are expected to be star forming.
@@
We find that on average, $\eta\approx\mathrm{const}$ is required to produce flattening in the luminosity–metallicity relation. Rey et al. (2025), for example, find that the inclusion of radiative feedback in addition to SN feedback in their EDGE simulations reduces $\eta$ and leads to higher metallicity in UFDs, making a better fit to luminosity–metallicity observations. In simulations with radiative feedback, the time-averaged $\eta$ is relatively constant in stellar mass for their sample of UFDs, with the $68\%$ interval between a few tenths to several hundred over the cosmological history of a UFD. This agrees with $\eta\approx\rm const$ as an explanation for the flattening of the $\mathrm{[Fe/H]-M_{V}}$ relation in the UFD regime, and broadly with our required range of $\eta_{\rm max}$ values.
Rather than regulating star formation with different modes of ISM heating, the outflow wind model in TNG50 determines wind injection velocity by the local dark matter velocity dispersion, subject to a minimum velocity requirement, therefore the winds launched at the injection scale reach a well defined maximum that leads to flattening in $\eta$ at this scale (Pillepich et al., 2018; Nelson et al., 2019b). However, as Nelson et al. (2019b) show, the emergent outflows at larger radii do not necessarily show the same flattening, subject to the radial velocity and distance used to define galactic outflow.
We note that our compilation of observed UFD mean metallicities are taken from spectroscopically derived values in the Local Volume Database (Pace, 2024). These are compiled from a range of studies, but predominantly originate from medium-resolution spectroscopic observations where metallicities are derived from the near-infrared calcium triplet lines (e.g. Li et al., 2017; Longeard et al., 2018; Jenkins et al., 2021; Bruce et al., 2023). Metallicity calibrations from these features are reliable down to $\mathrm{[Fe/H]}=-4.0$ with a minimum precision of $\sim 0.16$ dex (e.g. Carrera et al., 2013) that varies based on the S/N of the observed spectrum of each star. This metallicity calibration, and other typical spectroscopic calibrations (e.g., the Ca II K line; Beers et al. (1999)) or methods (i.e., high-resolution spectroscopy) do not impose an artificial metallicity floor at $\mathrm{[Fe/H]}\sim-3.0$. Accordingly, it is unlikely that the leveling of the metallicity-luminosity relation is due to an artifact in the derived metallicities.
@@ 5 Summary and Conclusions
    We examine the metallicity distribution of the IGM gas that would be accretable by UFD progenitor galaxies at $5<z<10$ in several state-of-the-art simulations of galaxy formation with different treatments of the gas thermodynamics and radiative transfer, star formation, Pop III modelling, and stellar feedback. We find that although the distributions are fairly broad, most of the available IGM gas has low metallicity with only a few per cent of the IGM available to UFDs enriched to $\mathrm{[Fe/H]}\geq-4$ (see Figure 2).